OpenStudy (anonymous):

soybean meal is 18% protein, cornmeal is 9% protein. How many pounds should be mixed of each to get 360 pounds of mixture that is 17% protein?

OpenStudy (anonymous):

I've seen this question 4 times on Openstudy :-P x + y = 360 0.18x + 0.9y = 0.17*360

OpenStudy (anonymous):

oops it's 0.09y Solve using gaussian elimination

OpenStudy (anonymous):

lol - must be a standard ques!! I am still confused

OpenStudy (anonymous):

I need hoe many pounds of cornmeal and how many pounds of soybean meal

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=rref%28 [[1%2C1%2C360]%2C[0.18%2C0.09%2C+0.17*360]]%29

OpenStudy (anonymous):

link is broken :-P

OpenStudy (anonymous):

\[\url{ http://www.wolframalpha.com/input/?i=rref%28 [[1%2C1%2C360]%2C[0.18%2C0.09%2C+0.17*360]]%29}\]

OpenStudy (anonymous):

still broken :-(

OpenStudy (anonymous):

so 320 of cornmeal and 40 of soybean????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you ROCK!!! Thanks

OpenStudy (anonymous):

How about checking out the oj prob?!?!?! :oP